1. Using the table from the last question, calculate the overall attack rate in persons who did not eat ice cream.

Ate ice cream Diarrhea 39 11 Total 52 15 Did not eat ice cream Diarrhea 14 Total 40 30 Ate pizza Did not eat pizza
Answer #1

Calculation of the overall attack rate in persons who did not eat ice cream:

Overall attack rate = Total number of new cases / Population X 100

Total no of new cases who did not eat ice cream = 14 + 9 = 23

Population = 40 + 30 = 70

Overall attack rate = 23 / 70 X 100

=32.85 percent.

= 33 percent.